mikewwe13:

Graph the system of equations on graph paper. {4x + 3y = 12 {2x + 3y = −6 Which statements are true about the solution to the system of equations? Select each correct answer. The x-coordinate of the solution is 8. The y-coordinate of the solution is 9. The y-coordinate of the solution is −8 . The x-coordinate of the solution is 9. The ordered pair that is the solution to the system lies in Quadrant I . The ordered pair that is the solution to the system lies in Quadrant IV .

Vocaloid:

the numbers work out well enough to find the x and y-intercepts for each graph instead of writing them in y = mx + b form let's start with 4x + 3y = 12 to find the x-intercept, let y = 0 and solve the equation for x

mikewwe13:

it's the first, fourth, and sixth

Vocaloid:

please follow my instructions you selected "The x-coordinate of the solution is 8." and "The x-coordinate of the solution is 9." which is impossible.

mikewwe13:

ok it's not them ok, cancel those out

1 attachment
mikewwe13:

both the equations is the 2nd and 5th ones

mikewwe13:

i looked over

Vocaloid:

please graph both lines.

mikewwe13:

1 attachment
Vocaloid:

goood now find: the x-coordinate the y-coordinate and the quadrant.

mikewwe13:

the x- coordinate is 9 the y-coordinate is -8 and the quadrant is IV

Vocaloid:

well done so statements 4, 3, and 6 = your answers

mikewwe13:

thank you so much, i really appreciate your help everytime

mikewwe13:

Use a graphing tool to solve the system. {4x − 15y = −43 {2x −3y = −2 Which ordered pair is the best estimate for the solution to the system? A. (5.5,4.3) B. (6.5,5.2) C. (4.8,5.2) D. (5.1,4.8)

Vocaloid:

re-write both equations in terms of y = mx + b and try to estimate where the graphs intersect as a hint, start by solving both equations for y

mikewwe13:

C or B

mikewwe13:

I pick B

Vocaloid:

good, b is correct

mikewwe13:

Graph the system of equations on graph paper. {6x + 4y = 24 {2x − 4y = −8 Which statements are true about the solution to the system of equations? Select each correct answer. The ordered pair that is the solution to the system lies in Quadrant I . The x-coordinate of the solution is 3. The x-coordinate of the solution is 2. The y-coordinate of the solution is −2 . The y-coordinate of the solution is 3. The ordered pair that is the solution to the system lies in Quadrant IV .

Vocaloid:

this one is pretty easy to graph, simply find the intercepts using the method I described earlier, or convert them both to y = mx + b form then find the solution and quadrant

mikewwe13:

1 attachment
Vocaloid:

check your graph again

Vocaloid:

one of the slopes needs to be positive

mikewwe13:

1 attachment
Vocaloid:

in this new graph both of your slopes are still negative

Vocaloid:

try solving both equations for y and re-write them in y = mx + b form and you will notice the mistake

mikewwe13:

but i already know the answers

Vocaloid:

it is asking you to graph the equation so you must provide an accurate graph.

mikewwe13:

i've seen my mistake, my apologizes

mikewwe13:

1 attachment
Vocaloid:

excellent, so find the x-coordinate y-coordinate and the quadrant

mikewwe13:

the x-coordinate is 2 the y-coordinate is 3 and the quadrant is I

mikewwe13:

AM I RIGHT ?

Vocaloid:

good, so statements 2, 4, 1 = your answer

mikewwe13:

so wait is x coordinate 2 or 3 and y is -2 ?

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mikewwe13 the x-coordinate is 2 the y-coordinate is 3 and the quadrant is I \(\color{#0cbb34}{\text{End of Quote}}\)